7. An oxygen tank in a laboratory has an internal pressure of 12 atm at 25°C. If the pressure inside the tank exceeds 25 atm, #8 please
Answer #1

Sol 8.

As the number of moles of nitrogen gas in the bubble is constant , so ,

P1 × V1 / T1 = P2 × V2 / T2  

where P1 = Pressure at depth of 105 feet = 4.15 atm

V1 = Volume of nitrogen gas bubble in the knee

= 0.0125 mL

T1 = Temperature at depth of 105 feet

= 10.5°C = 283.65 K

P2 = Pressure at surface = 0.998 atm

T2 = Temperature at surface = 29.5°C = 302.65 K

Therefore ,  

Volume of nitrogen gas bubble at the surface  

= V2

= P1 × V1 × T2 / ( P2 × T1 )

= 4.15 × 0.0125 × 302.65 / ( 0.998 × 283.65 )

= 0.0555 mL